OpenStudy (anonymous):

Evaluate the function f(x)=x^2-3x+7 Given value f(x+4)? Can you give me the answer to check if its right

OpenStudy (anonymous):

To solve this we need to substitute the given value into the original equation. In this case that amounts to replacing every x with x+4 and simplifying: \[f(x)=x^{2}-3x+7\] \[f(x+4)=(x+4)^{2}-3(x+4)+7\] \[f(x+4)=x^{2}+8x+16-3x-12+7\] \[f(x+4)=x^{2}+5x+11\]